OpenStudy (anonymous):

15. If the number of professors in a college is P and the number of students is S, and there are 20 times as many students as professors, write an algebraic equation that shows this relationship.

OpenStudy (anonymous):

20x + x = P + S 21x = P + S

OpenStudy (anonymous):

If there is one professor, there are 20 students. If there are two professors, there are 40 students. If there are five professors, there are 100 students. ... If there are P professors, there must be 20P students. So the number of students is 20P. So S = 20P. (Work your way up from easy numbers.)